OpenStudy (anonymous):

Perform the following multiplication. The answer should be in simplified form for radical expressions.

OpenStudy (anonymous):

\[\sqrt{10}(\sqrt{7}+\sqrt{10})\] I got \[\sqrt{170}\], but it says that's incorrect...

OpenStudy (anonymous):

Do I factor \[\sqrt{10}\]?

OpenStudy (anonymous):

I can't...

OpenStudy (sidsiddhartha):

its sqrt(170)+sqrt(10*10)=sqrt(170)+10

OpenStudy (anonymous):

Oooooh, cuz they're multiplied, so the radical is removed. got it! thanks !

OpenStudy (sidsiddhartha):

:)

OpenStudy (anonymous):

\[\sqrt{170}+10\]?

OpenStudy (sidsiddhartha):

yes

OpenStudy (anonymous):

It says it's not right :/

OpenStudy (anonymous):

Oooooh, it's just \[\sqrt{70}+10\]!